OpenStudy (king):

\[\log_{3} (\log_{2} x) + \log_{1/3} (\log_{1/2} y)]=1

OpenStudy (anonymous):

\[\log_{3} (\log_{2} x) + \log_{1/3} (\log_{1/2} y)=1\]

OpenStudy (king):

ya ure rite !!!!!!!now pls solve!

OpenStudy (anonymous):

you can't solve because you have two variables and one equation

OpenStudy (slaaibak):

Chuck norris can

OpenStudy (king):

oh sorry wait also given xy^2=4

OpenStudy (king):

\[xy ^{2}=4\]

OpenStudy (king):

now can u solve?

OpenStudy (slaaibak):

Yep

OpenStudy (anonymous):

\[\begin{cases}\log_{3} (\log_{2} x) + \log_{1/3} (\log_{1/2} y)=1\\xy ^{2}=4\end{cases}\]

OpenStudy (king):

so please solve.........

OpenStudy (slaaibak):

\[\log_3 (\log_2 x) - \log_3(-\log_2 y) = 1\] \[\log_3({ \log_2 x \over -\log_2 y }) = 1\] \[3 = { \log_2 x \over -\log_2 y }\] x=y^2/4 \[3 ={{ {\log_2{ y^2\over4}} } \over -\log_2 y}\] I'm not sure if I did everything correctly, but you can try to solve it further. If I made no mistakes, it's not too hard from here

OpenStudy (slaaibak):

Oops, there I see my mistake. x= 4/y^2 So just fix the last substitution

OpenStudy (king):

ummm.can u wait i shall try and get the answer and can u check if it is rite?

OpenStudy (slaaibak):

Yep

OpenStudy (king):

ummm...how do u get the answer??:(::(:(:(

OpenStudy (anonymous):

sorry but solving those logarithms is just waste of time, it's only technical thing, so you'd better show how you solve and then i can say what's wrong

OpenStudy (king):

but the thing is i dont know how to go on furtheer.......

OpenStudy (king):

ok guys i got this \[2-\log_{2} y ^{2}/\log_{2} y\]

OpenStudy (anonymous):

\[\log_nm^p=p\cdot\log_nm\] use this and stop spamming

OpenStudy (mertsj):